A $180,000 issue of seven-year bonds redeemable at par offers 6.24% coupons payable monthly. What is the premium or discount and the purchase price of the bonds to yield 9.6% compounded annually?

ANSWERS GIVEN

a) The DISCOUNT is $27,436.49

b) The purchase price of the bond is $152,563.51

**Note: would like to know how to get to these answers. Using a TI BA II PLUS. Please show Calculator solution (Easiest solution). Thank you *
Answer #1

YTM = 9.6% annually

APR(monthly) = 12[(1 + 0.096)1/12 - 1]

APR(monthly) = 9.202%

Par Value = $180,000

Coupon Payment = 180,000(0.0624/12) = $936

Time Period = 84 months

Calculating Bond Price,

Using TVM Calculation,

PV = [FV = 180,000, PMT = 936, N = 84, I = 0.09202/12]

PV = $152,563.51

Bond Price = $152,563.51

DIscount = 180,000 - 152,563.51

Discount = $27,436.49
Answer #2

Step-by-Step Calculation (Using TI BA II Plus):

1. Input Given Data:

  • Face Value (FV): $180,000

  • Coupon Rate: 6.24% annually → Monthly Coupon Payment:

    PMT=6.24%×$180,00012=$936

  • Time to Maturity (N): 7 years → 84 months (7 × 12)

  • Yield (YTM): 9.6% compounded annually → Monthly Equivalent Rate:

    (1+0.096)1/121=0.767% per month

2. Calculate the Purchase Price (PV):

  • Set Calculator to END Mode:

    • 2nd → PMT → 2nd → ENTER (to toggle BGN/END).

  • Input:

    • N = 84

    • I/Y = 0.767 (monthly rate)

    • PMT = 936

    • FV = 180,000

    • CPT → PV

Result:

PV=$152,563.51(Price paid)

3. Calculate the Discount:

Discount=Face ValuePurchase Price=$180,000$152,563.51=$27,436.49


Answer:
The bond has a discount of 27,436.49andapurchasepriceof

answered by: anonymous